OpenStudy (anonymous):

simplify the expression -6+i/-5+i

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

dies it really say "simplify"? instead of for example "write in standard form as \(a+bi\)"?

OpenStudy (misty1212):

multiply top and bottom by the "conjugate" of the denominator the conjugate of \(a+bi\) is (a-bi\) and this works because \[(a+bi)(a-bi)=a^2+b^2\] a real number

OpenStudy (misty1212):

i.e. compute \[\frac{-6+i}{-5+i}\times \frac{-5-i}{-5-i}\] the denominator will be \[56^2+1^2=26\] the numerator is whatever you get

OpenStudy (anonymous):

i dont understand how you got that do you mind explaining?

OpenStudy (anonymous):

the answer choices are a 31+i/26 b 29+i/26 c 31+i/24 d 31+11i/26